OpenStudy (anonymous):

How many milliliters of 0.30% sugar solution must you add to 75mL of 4% sugar solution to get a 0.50% sugar solution?

OpenStudy (kash_thesmartguy):

Please post it in Chemistry. This is math.

OpenStudy (anonymous):

but this is a math problem, I'm supposed to use rational functions to solve it

OpenStudy (mathstudent55):

75 ml of a 4% solution of sugar has 0.04(75 ml) amount of sugar. You are adding x ml of the 0.50% solution to it. How much sugar is there in x ml of 0.50% solution?

OpenStudy (anonymous):

.005x ml of sugar in the 0.50% solution

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

You need to add x amount of the 0.3% solution. In x ml of solution, the 0.3% solution has 0.003x amount of sugar. Ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now when you add the 75 ml of solution to x ml of 0.3% solution, the amount of sugar will be: \(3 + 0.003x\) Right?

OpenStudy (anonymous):

why is it 3?

OpenStudy (mathstudent55):

4% * 75 ml = 3 ml (this amount of sugar comes from the 4% solution.) 0.003x (this amount of sugar comes from the 0.3% solution) We add the sugar, and we get: \(3 + 0.003x\)

OpenStudy (anonymous):

okay I understand now

OpenStudy (anonymous):

4% * 75 ml = 3 ml (this amount of sugar comes from the 4% solution.) 0.003x (this amount of sugar comes from the 0.3% solution) We add the sugar, and we get: 3+0.003x yup he is correct next time post in chemestry though but i understand why you did it in math

OpenStudy (mathstudent55):

The concentration of sugar is the amount of sugar divided by the total volume. We divided the amount of sugar, 3 + 0.003x by x + 75 and we want that to be 0.5% \(\dfrac{3 + 0.003x}{75 + x} = 0.005\)

OpenStudy (mathstudent55):

This is a math problem that can very well be in a math book under rational equations.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The concentration of sugar is the amount of sugar divided by the total volume. We divided the amount of sugar, 3 + 0.003x by x + 75 and we want that to be 0.5% 3+0.003x/75+x =0.005 do you get it

OpenStudy (mathstudent55):

math deals with many situations, such as concentrations of solutions.

OpenStudy (anonymous):

yeah, I get that

OpenStudy (mathstudent55):

Now you solve the equation to find x, the volume of the 0.3% solution needed.

OpenStudy (anonymous):

I got x=1,312.5 Is that correct?

OpenStudy (mathstudent55):

Yes! Great job!

OpenStudy (anonymous):

oh! ok, thanks so much

OpenStudy (mathstudent55):

You're welcome.